The owner of the factory that sells a particular bottled fruit juice claimed that the

average capacity of their product is 250 ml. To test the claim, a consumer group gets

a sample of 100 such bottles, calculates the capacity of each bottle, and then finds

the mean capacity to be 248 ml. The standard deviation s is 5 ml. Use α = 0.03. Will

you reject or not to reject the claim.


1
Expert's answer
2022-05-30T05:39:54-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=250"

"H_1:\mu\not=250"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.03," "df=n-1=99" and the critical value for a two-tailed test is "t_c = 2.201819."

The rejection region for this two-tailed test is "R = \{t:|t|> 2.201819\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{248-250}{5/\sqrt{100}}=-4"

Since it is observed that "|t|=4> 2.201819=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=99" degrees of freedom, "t=-4" is "p=0.000122," and since "p=0.000122<0.03=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 250, at the "\alpha = 0.03" significance level.